OpenStudy (anonymous):

Hi I'm having problems solving probability. Can someone please help me? It is driving me crazy.

OpenStudy (dumbcow):

sure whats the question

OpenStudy (saifoo.khan):

in can try.

OpenStudy (anonymous):

K, thank u. One second...

OpenStudy (anonymous):

Ok, the question is: "What is the probability of answering at least 1 question out of 5 correctly?"

OpenStudy (anonymous):

..."Each question has 4 possible choices"

OpenStudy (saifoo.khan):

1/5

OpenStudy (anonymous):

I added what I thought were the probability of getting exactly 1, exactly 2 ... and so on to 5 up. But I did not get the right answer : (

OpenStudy (anonymous):

This is how I set it up: (.25*.75^4)+(.25^2*.75^3)+(.25^3*.75^2)+(.25^4*.75)+(.25^5)=

OpenStudy (dumbcow):

well its 1-prob of 0 out of 5 correctly since there are 4 choices, the probability of choosing the correct choice is 1/4 probability of choosing wrong answer is 3/4 probability of getting all questions wrong is (3/4)^5 So the answer is 1 - (3/4)^5

OpenStudy (anonymous):

1/4 = .25 = probability of getting an answer correct 3/4 = .75 = probability of getting an answer incorrect

OpenStudy (anonymous):

*reading your answer now...

OpenStudy (anonymous):

Try to find the probability of answering all the questions incorrectly, and lets call that P. The probability of answering at least one question correctly 1 - P, since it's the probability of not answering all questions incorrectly. The probability of answering one question incorrectly is: \[\frac{3}{4},\]since three out of four choices are wrong. Because there are five questions, you have to multiply this number by itself five times, so\[P = \left(\frac{3}{4}\right)^5\]and the probability of answering at least one question correctly is\[1-P = 1-\left(\frac{3}{4}\right)^5 = \frac{781}{1024} = 0.762695.\]

OpenStudy (anonymous):

ok, i think I understand. so does that mean that the probability of getting at lest 2 correct = 1 - (3/4)^5 ?

OpenStudy (anonymous):

Also, is it always a good idea to do "at least" problems as "1 - probability of the opposite"?

OpenStudy (dumbcow):

math_moron, you actually almost had it set up correctly you just forgot the coefficients, 5Cx where x is the number of correct questions

OpenStudy (dumbcow):

5(.25*.75^4)+10(.25^2*.75^3)+10(.25^3*.75^2)+5(.25^4*.75)+(.25^5)

OpenStudy (anonymous):

Sorry I'm just trying to wrap my head around what you did there

OpenStudy (anonymous):

No, the probability of getting at least one correct is\[1-\left(\frac{3}{4}\right)^5.\]The probability of getting at least two correct is\[P=1-\left(\frac{3}{4}\right)^5 - Q\]where Q is the probability of getting exactly one correct, which is\[Q = \binom{5}{1}\frac{1}{4}\left(\frac{3}{4}\right)^4\]that is, the probability of answering one correctly (1/4) times the probability of answering the other four incorrectly ((3/4)^4) times five which is the number of ways of choosing one question out of five, so\[P = 1 - \left(\frac{3}{4}\right)^5-5\frac{1}{4}\left(\frac{3}{4}\right)^4 = \frac{47}{128} = 0.3671875.\]

OpenStudy (anonymous):

ok, I think I'm starting to see it. Thank u all! I'm unfamiliar w/ how this site works as this is my first time. Is there a way for me to award you points for helping me? Also, is anyone aware of additional resources I can use for this topic?

OpenStudy (anonymous):

Click on the good answer button next to our answer. You can look up "binomial probability" for more information.

OpenStudy (anonymous):

Ok thank u for your help. I will look that up after I find this "good answer button" : )